OpenStudy (anonymous):

Please Help!!!!!! Solve the equations for the solutions, x and y . 2x + y = 3 x - 2y = -1 lines intersect at __, __

OpenStudy (anonymous):

(1,1)

OpenStudy (anonymous):

Solve the equations for the solutions, x and y . x - 2y = 3 5x + 3y = 2 The lines intersect at __, __

OpenStudy (anonymous):

(1,-1)

OpenStudy (anonymous):

Solve the equations for the solutions, x and y . x - y = 0 6x + 5y = -22 The lines intersect at __, __

OpenStudy (anonymous):

(-2,-2)

OpenStudy (anonymous):

Solve the equations for the solutions, x and y . 3x + 5y = 78 2x - y = 0 The lines intersect at

OpenStudy (anonymous):

@esamalaa

OpenStudy (anonymous):

(6,12)

OpenStudy (anonymous):

Solve the equations for the solutions, x and y . x + 5y = -2 2x + y = 5 The lines intersect at

OpenStudy (anonymous):

(3,-1)

OpenStudy (anonymous):

Solve the equations for the solutions, x and y . 2x + y = 1 9x + 3y = -3 The lines intersect at

OpenStudy (anonymous):

@esamalaa

OpenStudy (anonymous):

(-2,5)

OpenStudy (anonymous):

Solve the equations for the solutions, x and y . 5x + 2y = 6 3x + y = 4 The lines intersect at

OpenStudy (anonymous):

(2,-2)

OpenStudy (anonymous):

Solve the equations for the solutions, x and y . 3x + y = 4 x - 2y = 3 The lines intersect at

OpenStudy (anonymous):

(11/7,-5/7)

OpenStudy (anonymous):

Solve the equations for the solutions, x and y . 2x + 3y = 6 5x + 2y = 4 The lines intersect at

OpenStudy (anonymous):

@esamalaa